OpenStudy (bookworm14):

I got the last 3 problems correct and then my mind went blank ...help please ?!?!?!?! Daniel is 4 times as old as Jessica and is also 15 years older than Jessica. How old is Jessica?

OpenStudy (sleepyjess):

Daniel = d Jessica = j d=4j + 15

OpenStudy (blake57roger):

hi

OpenStudy (bookworm14):

I got that formula d=4j+15 , but all of a sudden my mind went blank and i forgot what to do x_x

OpenStudy (sleepyjess):

Is there any other information? Jessica could be any age with this information about the problem.

OpenStudy (sydthekid913):

I got the answer I think

OpenStudy (sleepyjess):

Or any answer choices?

OpenStudy (bookworm14):

no that is all the info given, the last three were the same and i figured those out

OpenStudy (sleepyjess):

Could you post one of the other 3 with the answer to see if that helps me figure this out?

OpenStudy (sydthekid913):

I got j=5 because if you multiply that by 4 you get twenty which is also. 15 years older

OpenStudy (bookworm14):

Sorry but I can't go back through the questions, this is on khanacademy.org so once you move on to the next question you cant go back

OpenStudy (sleepyjess):

Syd is right. The formula should be D+15=4J

OpenStudy (bookworm14):

@SydtheKid913 yes !!! that is how I worked out the other ones, thank you soooo much xD I have 1 left can yall help this one is much more complex ughh :(

OpenStudy (bookworm14):

so i had the info for the formula right, but the actual formula i was using was wrong darnnnn !!

OpenStudy (sleepyjess):

I will try. \(\huge\ddot\smile\) Go ahead and post it.

OpenStudy (sydthekid913):

Sure! Can you medal me? (Trying to boost smartscore):)

OpenStudy (sydthekid913):

O nevermind:)

OpenStudy (bookworm14):

Christopher is 2 times as old as Vanessa. 35 years ago, Christopher was 7 times as old as Vanessa. How old is Christopher now? p.s I could use some medals too if you don't mind ... p.s.s @sleepyjess how did you get that smiley face insert ? :O

OpenStudy (sleepyjess):

C=2V 35 years ago C-35=7V-35 Is what I am thinking for this one

OpenStudy (sleepyjess):

\(\huge\ddot\smile\ ) without spaces @Bookworm14

OpenStudy (sydthekid913):

\(\huge\ddot\smile\)

OpenStudy (sydthekid913):

Hey it worked!

OpenStudy (sleepyjess):

If you want to learn more I can teach you

OpenStudy (sydthekid913):

Whoa! That's AWESOME

OpenStudy (bookworm14):

okay how about we finish the math problem first and then we learn those cool tricks !?!?!?

OpenStudy (sleepyjess):

Ok \(\huge\color{#00bfff}{\ddot\smile}\)

OpenStudy (sydthekid913):

Cool

OpenStudy (sleepyjess):

What do you think of my equations @SydtheKid913

OpenStudy (sleepyjess):

C=2V 35 years ago C-35=7V-35

OpenStudy (sydthekid913):

I think they look hard :P

OpenStudy (sydthekid913):

Jk I'm still workin on it but I'm getting there

OpenStudy (bookworm14):

so here is what I got they might look a little different but im just guessing here C=2V (V-35)*7= C-35 Idk for sure though

OpenStudy (sleepyjess):

@ganeshie8

OpenStudy (sydthekid913):

C=2V 35 years ago C-35=7V-35

OpenStudy (sleepyjess):

Lets try a number for V and see what that does. Maybe 40?

OpenStudy (bookworm14):

@cody7744 @Cody17 @jakewoln @Faceomo @DangerousJesse @agent0smith @undeadknight26 okay sounds good @sleepyjess

OpenStudy (sydthekid913):

That's what I've been doing

OpenStudy (anonymous):

hi there

OpenStudy (sleepyjess):

C=2(40) c=80 80-35=(40-35)*7 45=5*7 45=35 V isnt 40

OpenStudy (bookworm14):

Maybe we should try a lower number? idk

OpenStudy (sydthekid913):

@haleyelizabeth2017, @iambatman, @ParthKohli

OpenStudy (bookworm14):

So I know that V has to be greater 35 thats for sure

OpenStudy (sydthekid913):

Yes

OpenStudy (sydthekid913):

So look for multiples of 7 greater than. 35

OpenStudy (sleepyjess):

C=2(37) C=74 74-35=(37-35)*7 39=2*7 39=14 not 37 c=2(42) c=84 84-35=(42-35)*7 49=7*7 49=49 YYYYYYAAAAAAAAYYYYYYY

OpenStudy (bookworm14):

No way I was just about to tell yall that I though the answer was 42 !!!!! now lets see if its right ...................

OpenStudy (sleepyjess):

I just tested it up there lol

OpenStudy (sydthekid913):

I got it!

OpenStudy (sydthekid913):

I got the answer!

OpenStudy (sleepyjess):

Christopher is 84 now.

OpenStudy (bookworm14):

what the heck i entered it and it was wrong, so i looked at how they worked it out and the answer isnt even a whole number??? im gonna post a pic of how they worked it out and see if it makes sense to yall ????

OpenStudy (sydthekid913):

C=84 V=42 35 years ago V=7 and C=49

OpenStudy (bookworm14):

wait no the correct answer was 84 i forgot to multiply by 2 x_x great now i have to do five more of these (facepalms)

OpenStudy (sleepyjess):

NOOOOOOO!!!!!!!! lol

OpenStudy (sydthekid913):

I just posted the answer you didn't have to get it wrong :/

OpenStudy (bookworm14):

Im doomed I should pay more attention ughhhhhh

OpenStudy (sleepyjess):

We will help you. Or at least I will. I don't know about @SydtheKid913

OpenStudy (bookworm14):

thank you @sleepyjess that means alot :') \(\huge\ddot\smile\)

OpenStudy (bookworm14):

okie dokie here we go .... Gabriela is 30 years younger than Omar. For the last four years, Omar and Gabriela have been friends. Eighteen years ago, Omar was 4 times as old as Gabriela. How old is Omar now?

OpenStudy (sydthekid913):

Hey!! @sleepyjess I'll help too!

OpenStudy (sleepyjess):

G=O-30 (G-18)*4=O-18

OpenStudy (sleepyjess):

I didn't know if you were still available so I didn't want to commit you to helping lol @SydtheKid913

OpenStudy (sydthekid913):

Lol XD

OpenStudy (sydthekid913):

Someone has got to help me with my algebra1b when were done here though

OpenStudy (bookworm14):

okay so i got the same formula lets start substituting #'s , in this one we know that Gabriella is gonna have to be older than 18 @SydtheKid913 if I have time I will do my best to help you :)

OpenStudy (sydthekid913):

Thanks! What math is this?

OpenStudy (sleepyjess):

Lets see if Omar is 60 G=60-30 G=30 (30-18)*4=60-18 12*4=42 48=42 darn almost Maybe he is 52? G=52-30 G=22 (22-18)*4=52-18 4*4=34 16=34 nope 56? G=56-30 G=26 (26-18)*4=56-18 8*4=38 32=38 grrrrr 58 it has to be G=58-30 G=28 (28-18)*4=58-18 10*4=40 40=40 YYYYYYYYAAAAAAAYYYYYYY

OpenStudy (sydthekid913):

Good job!

OpenStudy (bookworm14):

It is Algebra 1 and im so frustrated it just logged me out and then I clicked back on it and it gave me a dif problem >:( I can't stand technology sometimes !! @sleepyjess I was almost there I was working out if he was 55 lol

OpenStudy (sleepyjess):

I was just starting with a low and a high number and going from there

OpenStudy (bookworm14):

okay next problem: (thank you so much for helping meee ) :) this one seems a little easier: William is 4 times as old as Ashley and is also 27 years older than Ashley. How old is William?

OpenStudy (sleepyjess):

W=4A OR W-27=A

OpenStudy (sleepyjess):

We know that William has to be at least 27

OpenStudy (bookworm14):

yep

OpenStudy (sleepyjess):

Lets try A is 7 W=4(7) W=28 nope 9? W=4(9) W=36 36-27=9 WHOOOOOO HOOOOO

OpenStudy (sydthekid913):

Wow u r really good at these!

OpenStudy (sleepyjess):

just trial and error \(\huge\color{goldenrod}{\ddot\smile}\)

OpenStudy (bookworm14):

math is usually my best subject but i just can't seen to work these out as fast as you holy cow !!! so isn't 9 how old ashley is?

OpenStudy (sydthekid913):

Why are there so many people watching this question?

OpenStudy (sleepyjess):

yes ashley is 9 but you want the age of william which is 36

OpenStudy (bookworm14):

okay fwew thats what i thought thank you !!

OpenStudy (bookworm14):

it was CORRECT !! wooohoooo lets keep going thank you thank you thank you !!!

OpenStudy (sleepyjess):

YAY!!!!

OpenStudy (sleepyjess):

brb

OpenStudy (bookworm14):

Umaima is 26 years old and Michael is 4 years old. How many years will it take until Umaima is only 3 times as old as Michael? @sleepyjess k see un in a min.

OpenStudy (sydthekid913):

Hey I would love to help with the rest but I REALLY need help on my math so I'll post my question and come back when it's answered

OpenStudy (bookworm14):

ok thats fine !!

OpenStudy (sleepyjess):

Ok i back

OpenStudy (sleepyjess):

This one is really just trial and error. I don't know how to make a formula for this so lets just try numbers. I have already done up to Michael being 12 in my head.

OpenStudy (sydthekid913):

\(\huge\ddot\smile\)

OpenStudy (sleepyjess):

Wait let me try 11 m=11 11-4=7 years 26+7=33 11*3=33 YAY!

OpenStudy (bookworm14):

I think Michael would be 11 when Umaima was was 33 so that is what age they are when he is 3x as old, so 11-4= 7, so it took 7 years for umaima to be 3x as old as michael ,,, do you agree?

OpenStudy (bookworm14):

i submitted it and we ............ GOT IT RIGHT !!! WOOOO lol xD

OpenStudy (bookworm14):

ohhh no its another one of those hard ones xD (well kinda hard) : Stephanie is 24 years younger than Kevin. Six years ago, Kevin was 4 times as old as Stephanie. How old is Kevin now?

OpenStudy (bookworm14):

k-24=s k-6= (s-6) * 4 I think thats the right formula?

OpenStudy (sleepyjess):

S=K-24 S-6=(K-6)*4